Light travels through a medium with an index of refraction of 1.04.It comes to the surface of another medium at an angle of 83 degree. If the index of refraction for this medium is 2.1, what angle does the refracted light make with respect to the surface?
1
Expert's answer
2021-09-27T09:01:59-0400

"n_1 = 1.04;\theta_1=83\degree"

"n_2 = 2.1;"

"\text{by Snell's Law:}"

"n_1\sin\theta_1=n_2\sin\theta_2"

"sin\theta_2=\frac{n_1\sin\theta_1}{n_2}=\frac{1.04\sin83\degree}{2.1}\approx0.4915"

"\theta_2 = 29.44\degree"


"\text{Answer: }\theta_2 = 29.44\degree"